Question 8 (1 point) d In a certain high school, the probability that a student drops out is 0.09, and the probability that a dropout gets a high-school equivalency diploma (GED) is 0.27. What is Write only a number as your answer. Round to four decimal places (for example: 0.2711). Do not write as a percentage.
Answer #1

6) Mean = (0*27 + 1*170 + 2*304 + 3*272 + 4*19 + 5*10) / (27 + 170 + 304 + 272 + 19 + 10) = 2.14 (ans)

8) P(student drops out) = 0.09

P(student gets GED | drops out) = 0.27

P(students gets GED) = P(student gets GED | drops out) * P(student drops out) = 0.27 * 0.09 = 0.0243 (ans)